stud
bud
Sign up
Log in
Home
Library
Explore
SAT
More
Statistics 1: Probability 1 Flashcards | StudBud
Explore
/
Statistics 1: Probability 1
Statistics 1: Probability 1
19 terms
·
StudBud Administration
·
32 plays
Imported from external platform
Flashcards
19 cards
F
Practice quiz
P
Host live
L
Questions
19
Answers
1
What is the IQR of this box and whisker plot?
85
70
45
15
2
At which value is the median?
5
7
4
8
3
How many total possible outcomes does this tree diagram have?
18
3
4
9
4
Use the multiplication rule!
.0046
.0833
1.00
.0278
5
What's the definition of independent events?
Two events can happen at the same time
One event doesn't influence the next
Two events can't happen at the same time
The outcome of one event influences the outcome of the next
6
What is the range of this box and whisker plot?
100
35
80
65
7
The multiplication rule is...
the same for independent and dependent events
none of the above
different for independent and dependent events
it depends
8
Which value could not represent a probability?
2/5
10%
0.00454
1.0035
9
What is Q3 of the box and whisker plot?
23
10
24
16
10
What is the probability of picking a male?
21/240
60/240
120/240
39/240
11
What is probability?
difference with how something will always happen verse never
The measure of likely hood to get something.
The law that describes how to get a better likely hood
The explantation of how to get a better chance at something.
12
How many outcomes does this tree diagram have?
10
4
6
8
13
Select the term or phrase which is NOT typically associated with the Addition Rule.
in the same trial
or
independent events
mutually exclusive events
14
of A and B which is the mean and which is the median?
A- Median B- Mean
A- Mean B- Mode
A-Mean B-Mean
A- Mean B- Median
15
According to the picture, determine the probability of getting a
green
ball.
1/4
0
1
5/12
16
What does a box and whisker plot contain?
Median
Mode
Outlier
Mean
17
Which formula is the General Addition Rule?
P(A and B) = 0
P(A or B) = P(A) + P(B)
P(not A) = 1 - P(A)
P(A or B) = P(A) + P(B) - P(A and B)
18
If A and B are independent events, then
P(AB) = P(A)P(B)
P(A-B) = P(A) - P(B)
P(A|B) = P(A)/P(B)
P(A+B) = P(A) + P(B)
19
What is conditional probability ?
(m)/(m+n)
a controlled trial
probability of event B given event A
probability of an event happening